# Quantum optical coherence theory based on Feynman’s path integral

### v1

@@ Abstract
The first, second and higher-order interference of classical and nonclassical light is interpreted in quantum optical coherence theory based on Feynman’s path integral. Compared to the existed classical optical coherence theory based on Maxwell’s electromagnetic theory and Glauber’s quantum optical coherence theory based on wave mechanics formulation of quantum mechanics, quantum optical coherence theory based on Feynman’s path integral provides a new tool to study optical coherence and has the advantage of understanding the connection between the physics and mathematical calculations. For instance, based on the results of transient first-order interference of two independent light beams, it is predicted that the well-accepted electric field model for thermal light introduced by many optical textbooks may not be accurate. The physics of two-photon bunching of thermal light and Hong-Ou-Mandel dip of entangled photon pairs is the same, which can be interpreted by constructive and destructive two-photon interference, respectively. The second-order interference of two independent light beams are introduced in detail. Interpreting the interference of light in quantum optical coherence theory based on Feynman’s path integral is helpful to understand the physics of light, and may eventually lead to the answer of the question puzzling us for a long time: what is photon?
@@ I Introduction
Ever since the interference of light introduced by Young in 1801 Young (1802), optical coherence has been one of the most important topics in the development of physics. With the contributions of generations of scientists, there are two main theories of optical coherence Born and Wolf (2013); Mandel and Wolf (1995). One is classical optical coherence theory based on Maxwell’s electromagnetic theory Born and Wolf (2013). The other one is Glauber’s quantum optical coherence theory based on wave mechanics formulation of quantum mechanics Glauber (1963a, c). It was proved by Glauber and Sudarshan independently that classical and quantum optical coherence theories are mathematically equivalent for classical light Sudarshan (1963); Glauber (1963b). However, classical optical coherence theory can not interpret non-classical phenomena such as photoelectric effect Hertz (1887); Einstein (1905), photon antibuching Miller and Mishkin (1967); Stoler (1974); Kimble et al. (1977), Bell inequality violation Bell (1964, 1966) and so on. On the other hand, quantum optical coherence theory can interpret both classical and non-classical phenomena.
@@
Since Glauber’s quantum optical coherence theory can explain all the optical coherence phenomena and be a great success Glauber (2006), what is the reason to introduce another quantum optical coherence theory based on Feynman’s path integral? This is the question we asked ourselves since the first day we employed Feynman’s path integral to explain the subwavelength interference of light Liu and Zhang (2010). Similar question bothered Young and Feynman Young (1802); Feynman (1948), too. In the beginning of Young’s paper on the theory of light and colours Young (1802), he explained why he employed wave theory to interpret light and colours: Although the invention of plausible hypotheses, independent of any connection with experimental observations, can be of very little use in the promotion of natural knowledge; yet the discovery of simple and uniform principles, by which a great number of apparently heterogeneous phenomena are reduced to coherent and universal laws, must ever be allowed to be of considerable importance towards the improvement of the human intellect.
@@
When Feynman introduced path integral formulation of quantum mechanics, he also faced the same question. There had been wave mechanics and matrix mechanics formulations of quantum mechanics. What is the reason to introduce a third and equivalent formulation of quantum mechanics. Feynman’s answer to this question is Feynman (1948): The formulation is mathematically equivalent to the more usual formulations. There are, therefore, no fundamentally new results. However, there is a pleasure in recognizing old things from a new point of view. Also, there are problems for which the new point of view offers a distinct advantage.
@@
The answers offered by Young and Feynman are also the reasons why we introduced quantum optical coherence theory based on Feynman’s path integral besides Glauber’s quantum optical coherence theory. As introduced in Sect. III, quantum optical coherence theory based on Feynman’s path integral can indeed offer deeper insight about the first- and second-order interference of two independent light beams, which is helpful to understand the physics of coherence and the concept of photon.
@@
The remaining parts of the paper are organized as follows. Basic concepts in quantum optical coherence theory based on Feynman’s path integral are introduced in Sect. II. The first- and second-order interference of light are introduced in Sects. III and IV, respectively. Generalizing the method developed in Sects. III and IV to the third- and higher-order interference of light can be found in Sect. V. Section VI introduces the quantum atomic coherence theory based on Feynman’s path integral by analogy of the optical one. Section VII summarizes the conclusions.
@@ II.1 History of quantum optical coherence theory based on Feynman’s path integral
The history of optical coherence began in the 17th century when Grimaldi discovered diffraction of light Hecht (2012). With the contributions from Young Young (1802, 1804), Fresnel Fresnel (1819), and many other scientists Hecht (2012); Born and Wolf (2013), classical optical coherence theory has been established based on Maxwell’s electromagnetic theory Maxwell (1865). We will not introduce the history of optical coherence since there are many good reviews about this topic Born and Wolf (2013); Hecht (2012); Mandel and Wolf (1965).
@@
The history of quantum optical coherence theory dates back to Dirac and Feynman. In his famous book on quantum mechanics, Dirac discussed the first-order interference of light with the concept of photon and concluded “Each photon then interferes only with itself. Interferences between two different photons never occurs Dirac (1930).” Feynman employed path integral to discuss Young’s double-slit interference of light in his lectures Feynman et al. (2011). All the optical coherence discussions were limited to the first-order before 1956. Both quantum and classical optical coherence theories can explain the first-order interference of light. The need for quantum optical coherence theory was not so urgent. In 1956, Hanbury Brown and Twiss found that photons emitted by thermal light source are not independent Brown and Twiss (1956a, b), which is known as two-photon bunching of thermal light. People began to realize that there are second- and higher-order interference of light besides the first-order. In 1963, Glauber published a series of papers explaining optical coherence in quantum theory Glauber (1963a, c, b), which laid the foundation of quantum optical coherence theory. Glauber was awarded 2005 Nobel physics prize for his contribution to quantum theory of optical coherence Glauber (2006). Glauber’s quantum optical coherence theory was employed by most scientists and introduced in many quantum optics books Mandel and Wolf (1995); Scully and Zubairy (1997); Loudon (2000); Shih (2020).
@@
Glauber’s quantum optical coherence theory is based on wave mechanics formulation of quantum mechanics Schrödinger (1926). There are two more equivalent formulations of quantum mechanics called matrix mechanics Heisenberg (1925, 1949) and Feynman’s path integral Feynman (1948); Feynman et al. (2010). In principle, one can establish quantum optical coherence theory based on matrix mechanics and Feynman’s path integral formulations, too. In the following parts of this paper, we will introduce quantum optical coherence theory based on Feynman’s path integral.
@@
Quantum optical coherence theory based on Feynman’s path integral was first introduced by Feynman himself. In his lectures on physics, Feynman examined Young’s double-slit interference of electrons, which is similar as the Young’s double-slit interference of light Feynman et al. (2011). Shortly after Hanbury Brown and Twiss’s experiments Brown and Twiss (1956a, b), Fano employed Feynman’s path integral to interpret two-photon bunching of thermal light Fano (1961). As shown in Fig. 1, there are two different ways for two photons emitted by atoms a and b and absorbed by atoms c and d. One is photon emitted by atom a is absorbed by atom c and photon emitted by atom b is absorbed by atom d. The second ways is emitted by atom a is absorbed by atom d and photon emitted by atom b is absorbed by atom c. Two-photon bunching of thermal light is interpreted as the superposition of these two different two-photon probability amplitudes corresponding these two different ways, which is later named as two-photon interference Shih (2020). In his lecture on quantum electrodynamics, Feynman also employed similar idea to interpret two-photon bunching of thermal light Feynman (2006).
@@
Figure 1: Interpretation of two-photon bunching of thermal light in Feynman’s path integral. From Fano (1961)
@@
As mentioned before, Glauber’s quantum optical coherence theory was employed to study coherence in most researches in quantum optics. There are limited number of studies employing Feynman’s path integral to study optical coherence. Hillery and Zubairy employed Feynman’s path integral to discuss some problems in nonlinear optics and calculated the path-integral representation for the coherent-state propagator Hillery and Zubairy (1982). Yabuki employed Feynman’s path integral to discuss the nonlinear contribution in Young’s double-slit interferometer of light Yabuki (1986). Barut and Basri discussed single-photon interference in single- and double-slit interferometers with Feynman’s path integral Barut and Basri (1992). Braidotti et. al. employed Feynman’s path integral to discuss non-paraxial optics problem Braidotti et al. (2018). Wallner employed Feynman’s path integral to discuss the interference of entangled photon pairs Wallner (2021).
@@
We have employed employed Feynman’s path integral to discuss the first- and second-order interference of light since 2010 Liu and Zhang (2010, 2011); Liu et al. (2013, 2014b, 2014a, 2015b, 2015a); Bai et al. (2016); Zhou et al. (2017, 2019a); Luo et al. (2021a, b). Some interesting conclusions have been drawn. For instance, it is predicted that there is no transient first-order interference pattern by superposing two independent thermal light beams Bai et al. (2016), which is different from the observed transient first-order interference pattern by superposing two independent laser light beams Javan et al. (1962); Magyar and Mandel (1963); Lipsett and Mandel (1963). This distinction leads to an interesting conclusion that electric field models for laser and classical light within the coherence time should be different, which is contrary to the usual electric model introduced in classical optics textbooks Born and Wolf (2013). Detail discussions about this problem can be found in Sect. III.
@@
The researches on optical coherence with Feynman’s path integral did by other groups Fano (1961); Feynman et al. (2011); Feynman (2006); Hillery and Zubairy (1982); Yabuki (1986); Barut and Basri (1992); Braidotti et al. (2018); Wallner (2021) and our group Liu and Zhang (2010, 2011); Liu et al. (2013, 2014b, 2014a, 2015b, 2015a); Bai et al. (2016); Zhou et al. (2017, 2019a); Luo et al. (2021a, b) are all about special interference phenomena of light. No systematic study about quantum optical coherence theory based on Feynman’s path integral is reported. This paper will present a comprehensive introduction about quantum optical coherence theory based on Feynman’s path integral by re-visiting the first-, second-, third- and higher-order interference of light.
@@ II.2 Principles of quantum optical coherence theory based on Feynman’s path integral
In order to calculate the probability of an event with the above assumptions, there are two premises need to be figured out. The first one is how to calculate the probability amplitude $\phi$ for each alternative way. The second one is how to decide whether two different alternative ways for an event are distinguishable or not.
@@ II.2.1 Probability amplitude for a path
The probability amplitude for a path is first given by Dirac by analogy of classical theory of dynamics Dirac (1930). Dirac assumed that the solution of schrödinger wave equation is as follows,
@@
Based on the above assumption, Dirac concluded that the probability amplitude can be expressed as a result of integral over all the possible paths
@@
where $t_{j}=t_{a}+(j-1)\times(t_{b}-t_{a})/m$ ($j=1,2,...,m$) and $q_{t_{j}}$ is the corresponding canonical coordinate at time $t_{j}$. Equation (2) is in fact path integral. The same equation can be found in Feynman’s book Feynman et al. (2010). The difference between Dirac and Feynma’s work on path integral is that Dirac gave the probability amplitude by analogy of classical dynamics and he used the word, analogous to, while Feynman showed that the probability amplitude in quantum mechanics is equal to Eq. (1) Feynman et al. (2010).
The probability amplitude for photon, usually called photon’s Feynman propagator, has simple form in Feynman’s Gauge in momentum space Greiner and Reinhardt (2013)
@@
where $g^{\mu\nu}$ is metric tensor in four-dimensional Minkowski spacetime, $k$ is four-dimension wave vector, $(k_{0},k_{1},k_{2},k_{3})$, $k_{0}=\omega$, $(k_{1},k_{2},k_{3})$ are three components of usual wave vector, $k^{2}=k^{\mu}k_{\mu}=k_{0}^{2}-k_{1}^{2}-k_{2}^{2}-k_{3}^{2}$, $\epsilon$ is an infinitesimal quantity introduced to ensure convergence of the integral.
@@
where $x$ and $y$ are four-dimensional Minkowski coordinate vector, $x=(x_{0},x_{1},x_{2},x_{3})$, $x_{0}$ is time $t$, $(x_{1},x_{2},x_{3})$ is the three components of usual coordinate vector $\vec{x}$. Equation (4) can be further simplified as Greiner and Reinhardt (2013)
$D_{F}(x;y)=\int\frac{d^{3}k}{(2\pi)^{3}}\frac{i}{(2\omega_{k})^{2}}e^{-i[\omega(t_{y}-t_{x})-\vec{k}\cdot(\vec{y}-\vec{x})]},$ (5)
@@
Comparing photon’s Feynman propagator Eq. (5) with Green function in classical optics Born and Wolf (2013)
@@
these two functions have similar expression with different contribution for each momentum component. The reason is that all the paths are taken into account in Eq. (5) with equal amplitude, while all the paths are taken into account in Eq. (6) with unequal amplitudes.
@@
where one dimension $x$ is assumed for simplicity. The particle’s Feynman propagator can be expressed as
@@
where $x_{a}$ and $x_{b}$ are the starting and ending positions of the particle, respectively, $S_{cl}(a,b)$ is the classical action of the particle between $x_{a}$ and $x_{b}$, $F(t_{a},t_{b})$ is a function of the starting and ending time. Equation (8) means that the probability amplitude of particle going from $x_{a}$ to $x_{b}$ can be related to the classical propagation function with an extra function only dependent on the starting and ending time. The Lagrangian of free photon has the form of Eq. (7) Srednicki (2007). Equation (8) should be valid for photons. Classical Green function can be used as photon’s Feynman propagator in certain cases.

### v2

@@ Abstract
Compared to classical optical coherence theory based on Maxwell’s electromagnetic theory and Glauber’s quantum optical coherence theory based on matrix mechanics formulation of quantum mechanics, quantum optical coherence theory based on Feynman’s path integral formulation of quantum mechanics provides a novel tool to study optical coherence. It has the advantage of understanding the connection between mathematical calculations and physical interpretations better. Quantum optical coherence theory based on Feynman’s path integral is introduced and reviewed in this paper. Based on the results of transient first-order interference of two independent light beams, it is predicted that the classical model for electric field of thermal light introduced by classical optical textbooks may not be accurate. The physics of two-photon bunching of thermal light and Hong-Ou-Mandel dip of entangled photon pairs is the same, which can be interpreted by constructive and destructive two-photon interference, respectively. Quantum optical coherence theory based on Feynman’s path integral is helpful to understand the coherence properties of light, which may eventually lead us to the answer of the question: what is a photon?
@@ I Introduction
Ever since the interference of light was introduced by Young in 1801 Young (1802), optical coherence has been one of the most important topics in the development of physics. With contributions from generations of scientists, two main optical coherence theories have been established Born and Wolf (2013); Mandel and Wolf (1995). One is classical optical coherence theory based on Maxwell’s electromagnetic theory Born and Wolf (2013). The other one is Glauber’s quantum optical coherence theory based on matrix mechanics formulation of quantum mechanics Glauber (1963a, c). It was proved by Glauber and Sudarshan independently that classical and quantum optical coherence theories are mathematically equivalent for classical light Sudarshan (1963); Glauber (1963b). However, classical optical coherence theory can not interpret non-classical phenomena such as photoelectric effect Hertz (1887); Einstein (1905), photon antibuching Miller and Mishkin (1967); Stoler (1974); Kimble et al. (1977), Bell inequality violation Bell (1964, 1966) and so on. On the other hand, Glauber’s quantum optical coherence theory can interpret both classical and non-classical phenomena.
@@
As Glauber’s quantum optical coherence theory can explain all the optical coherence phenomena Glauber (2006), why introduce another quantum optical coherence theory based on Feynman’s path integral? This is the question we asked ourselves when we employed Feynman’s path integral to explain the subwavelength interference of light Liu and Zhang (2010). Similar question bothered Young Young (1802) and Feynman Feynman (1948), too. At the beginning of Young’s paper on the theory of light and colours Young (1802), he explained why he employed wave theory: Although the invention of plausible hypotheses, independent of any connection with experimental observations, can be of very little use in the promotion of natural knowledge; yet the discovery of simple and uniform principles, by which a great number of apparently heterogeneous phenomena are reduced to coherent and universal laws, must ever be allowed to be of considerable importance towards the improvement of the human intellect.
@@
When Feynman introduced path integral formulation of quantum mechanics, he also faced similar question. There were wave mechanics and matrix mechanics formulations of quantum mechanics. Why introduce a third and equivalent formulation of quantum mechanics? Feynman’s answer to the question is Feynman (1948): The formulation is mathematically equivalent to the more usual formulations. There are, therefore, no fundamentally new results. However, there is a pleasure in recognizing old things from a new point of view. Also, there are problems for which the new point of view offers a distinct advantage.
@@
The answers given by Young and Feynman are also the reasons why we introduced quantum optical coherence theory based on Feynman’s path integral. As introduced in Sect. III, quantum optical coherence theory based on Feynman’s path integral can indeed offer deeper insight about the physics of the first- and second-order interference of two independent light beams, which is essential to understand the nature of light.
@@
The remaining parts of the paper are organized as follows. A short history of quantum optical coherence theory based on Feynman’s path integral and some basic concepts are introduced in Sect. II. Applying path integral to interpret the first- and second-order interference of light are in Sects. III and IV, respectively. Generalizing the method developed in Sects. III and IV to the third- and higher-order interference of light can be found in Sect. V. Section VI introduces quantum atomic coherence theory based on Feynman’s path integral by analogy of quantum optical coherence theory based on Feynman’s path integral. Our conclusions are summarized in Sect. VII.
@@ II.1 History of quantum optical coherence theory based on Feynman’s path integral
The history of classical optical coherence theory began in the 17th century when Grimaldi discovered diffraction of light Hecht (2012). With the contributions from Young Young (1802, 1804), Fresnel Fresnel (1821), and many other scientists Hecht (2012); Born and Wolf (2013), classical optical coherence theory has been established based on Maxwell’s electromagnetic theory Maxwell (1865). The detail history of classical optical coherence can be found in Refs. Born and Wolf (2013); Hecht (2012); Mandel and Wolf (1965).
@@
The history of quantum optical coherence theory dates back to Dirac and Feynman Dirac (1930); Feynman et al. (2011). In his famous book on quantum mechanics, Dirac discussed the first-order interference of light with the concept of photon and concluded that Each photon then interferes only with itself. Interferences between two different photons never occurs Dirac (1930). Feynman employed path integral to discuss the first-order interference of electrons in a Young’s double-slit interferometer in his lectures Feynman et al. (2011). All the discussions about optical coherence were limited to the first-order before 1956. Both quantum and classical optical coherence theories can explain the first-order interference of light. The need for quantum optical coherence theory was not so urgent then. In 1956, Hanbury Brown and Twiss found that photons emitted by thermal light source are not independent Brown and Twiss (1956a, b), which is known as two-photon bunching of thermal light. People began to realize that there are second- and higher-order interference of light besides the first-order interference. In 1963, Glauber published a series of papers explaining optical coherence in quantum theory Glauber (1963a, c, b), which laid the foundation of quantum optical coherence theory. Glauber was awarded 2005 Nobel physics prize for his contribution to quantum theory of optical coherence Glauber (2006). Glauber’s quantum optical coherence theory was employed by most scientists in quantum optics and introduced in many quantum optics textbooks Mandel and Wolf (1995); Scully and Zubairy (1997); Loudon (2000); Shih (2020).
@@
Glauber’s quantum optical coherence theory is based on matrix mechanics formulation of quantum mechanics Heisenberg (1925, 1949). There are two more equivalent formulations of quantum mechanics called wave mechanics Schrödinger (1926) and Feynman’s path integral Feynman (1948); Feynman et al. (2010). In principle, one can establish quantum optical coherence theory based on wave mechanics and Feynman’s path integral formulations, too. In the following parts, we will introduce quantum optical coherence theory based on Feynman’s path integral.
@@
Quantum optical coherence theory based on Feynman’s path integral was first introduced by Feynman himself. In his lectures on physics, Feynman examined Young’s double-slit interference of electrons, which is similar as the Young’s double-slit interference of photons Feynman et al. (2011). Shortly after Hanbury Brown and Twiss’s experiments Brown and Twiss (1956a, b), Fano employed Feynman’s path integral to interpret two-photon bunching of thermal light Fano (1961). As shown in Fig. 1, there are two different and indistinguishable paths for two photons emitted by atoms a and b to be absorbed by atoms c and d. The first path is the photon emitted by atom a is absorbed by atom c and the photon emitted by atom b is absorbed by atom d. The second path is the photon emitted by atom a is absorbed by atom d and the photon emitted by atom b is absorbed by atom c. Two-photon bunching of thermal light is interpreted as the superposition of two different two-photon probability amplitudes corresponding to these two different paths, which is later named as two-photon interference Shih (2020). In his lecture on quantum electrodynamics, Feynman also employed similar idea to interpret two-photon bunching of thermal light Feynman (2006).
@@
Figure 1: Interpretation of two-photon bunching of thermal light in Feynman’s path integral Fano (1961). Atoms a and b are two atoms that can emit a photon each. Atoms c and d are two atom that can absorb a photon each. $u$ and $v$ are two paths.
@@
As mentioned before, Glauber’s quantum optical coherence theory was employed to study optical coherence in most researches in quantum optics. There are limited number of researches employing Feynman’s path integral to study optical coherence. Hillery and Zubairy employed Feynman’s path integral to discuss some problems in nonlinear optics and calculated the path integral representation for coherent state propagator Hillery and Zubairy (1982). Yabuki employed Feynman’s path integral to discuss the nonlinear contribution in the first-order interference of light in a Young’s double-slit interferometer Yabuki (1986). Barut and Basri discussed single-photon interference in the single- and double-slit interferometers with Feynman’s path integral Barut and Basri (1992). Braidotti et. al. employed Feynman’s path integral to discuss some non-paraxial optics problems Braidotti et al. (2018). Stöhr showed that Feynman’s path integral reduces to the conventional wave formalism of light diffraction in the lowest order Stöhr (2019). Wallner employed Feynman’s path integral to discuss the interference of entangled photon pairs Wallner (2021). Wen et al. experimentally measured Feynman’s propagator for single photons Wen et al. (2023).
@@
We have employed Feynman’s path integral to discuss the first- and second-order interference of light since 2010 and obtained some interesting conclusions Liu and Zhang (2010, 2011); Liu et al. (2013, 2014b, 2014a, 2015b, 2015a); Bai et al. (2016); Zhou et al. (2017, 2019a); Luo et al. (2021a, b). For instance, it is predicted that there is no transient first-order interference pattern by superposing two independent thermal light beams Bai et al. (2016), which is different from the one of superposing two independent laser light beams Javan et al. (1962); Magyar and Mandel (1963); Lipsett and Mandel (1963). The difference leads to an interesting conclusion that classical electric field models for laser and thermal light within the coherence time should be different, which is contrary to the well-accepted electric model introduced in classical optics textbooks Born and Wolf (2013). Detail discussions about this problem can be found in Sect. III.
@@
The researches on optical coherence with Feynman’s path integral performed by other groups Fano (1961); Feynman et al. (2011); Feynman (2006); Hillery and Zubairy (1982); Yabuki (1986); Barut and Basri (1992); Braidotti et al. (2018); Wallner (2021) and our group Liu and Zhang (2010, 2011); Liu et al. (2013, 2014b, 2014a, 2015b, 2015a); Bai et al. (2016); Zhou et al. (2017, 2019a); Luo et al. (2021a, b) are all about certain specified interference phenomena of light. No systematic study about quantum optical coherence theory based on Feynman’s path integral is reported. This paper presents a systematic introduction and review about quantum optical coherence theory based on Feynman’s path integral.
@@ II.2 Principles of quantum optical coherence theory based on Feynman’s path integral
There are two premises for calculating the probability of an event with the above assumptions. The first premise is how to calculate the probability amplitude, $\phi$, for each path. The second premise is how to decide two different alternatives for an event are distinguishable or not.
@@ II.2.1 Probability amplitude for a path
The probability amplitude for a path is first given by Dirac by analogy of action in classical dynamics theory Dirac (1930). Dirac assumed that the solution of schrödinger wave equation is
@@
Based on the above assumption, Dirac concluded that the probability amplitude of the path can be expressed as the integral of all the possible paths
@@
where $t_{j}=t_{a}+(j-1)\times(t_{b}-t_{a})/m$ ($j=1,2,...,m$) and $q_{t_{j}}$ is the corresponding canonical coordinate at time $t_{j}$. Equation (2) is in fact the key result for path integral. The exactly same equation can be found in Feynman’s book Feynman et al. (2010). The difference between Dirac and Feynma’s work on path integral is that Dirac gave the probability amplitude by analogy of classical dynamics and he used the word, analogous to, while Feynman proved that the probability amplitude in quantum mechanics is equal to Eq. (1) Feynman et al. (2010).
The probability amplitude for photons, usually called photon’s Feynman propagator, has a simple form in momentum space in Feynman’s Gauge Greiner and Reinhardt (2013)
@@
where $g^{\mu\nu}$ is the metric tensor in four-dimensional Minkowski spacetime, $k$ is a four-dimension wave vector, $(k_{0},k_{1},k_{2},k_{3})$, $k_{0}=\omega$, $(k_{1},k_{2},k_{3})$ are three components of usual wave vector, $k^{2}=k^{\mu}k_{\mu}=k_{0}^{2}-k_{1}^{2}-k_{2}^{2}-k_{3}^{2}$, $\epsilon$ is an infinitesimal quantity introduced to ensure convergence of the integral.
@@
where $x$ and $y$ are four-dimensional Minkowski coordinate vectors, $x=(x_{0},x_{1},x_{2},x_{3})$, $x_{0}$ is time $t$, $(x_{1},x_{2},x_{3})$ is the three components of the usual coordinate vector $\vec{x}$. Equation (4) can be further simplified as Greiner and Reinhardt (2013)
$D_{F}(x;y)=\int\frac{d^{3}k}{(2\pi)^{3}}\frac{i}{(2\omega)^{2}}e^{-i[\omega(t_{y}-t_{x})-\vec{k}\cdot(\vec{y}-\vec{x})]},$ (5)
@@
Comparing photon’s Feynman propagator, Eq. (5), with Green function in classical optics Born and Wolf (2013)
@@
it is easy to find out that these two functions have similar expressions with different contributions for each momentum component. The reason is that all the paths are taken into account in Eq. (5) with equal amplitude, while all the paths are taken into account in Eq. (6) with unequal amplitudes.
@@
the particle’s Feynman propagator can be expressed as
@@
where one dimension case is assumed for simplicity, $x_{a}$ and $x_{b}$ are the starting and ending positions of the particle, respectively, $S_{cl}(a,b)$ is the classical action of the particle, $F(t_{a},t_{b})$ is a function of the starting and ending time. Equation (8) means that the particle’s Feynman propagator going from $x_{a}$ to $x_{b}$ can be related to the classical propagation function with an extra function which is only dependent on the starting and ending time. The Lagrangian of free photons has the form of Eq. (7) Srednicki (2007). Equation (8) should be valid for photons. Classical Green function can be used as photon’s Feynman propagator in certain cases.